# Incoherent Approximation of Leakage in Quantum Error Correction

### v1

@@ Appendix E Logical error rate
Appendix F Additional RPA simulation data
Here we provide supplemental plots and data pertaining to the validation of the Random Phase Approximation, Sect. III.1. First, we show that the detection event statistics relating to the main text ‘thermal heating model’, Fig. 1, are consistent in the model with and without the RPA. A detection event is defined as a particular stabilizer value flipping from one round to the next (due to a physical error) chen_exponential_2021 . We compute the detection event fraction (DEF), which is the probability a given measure qubit (stabilizer) registers a detection event, in a given round. It is more instructive to plot the added DEF due to leakage, obtained by considering the difference in DEFs between simulations with and without leakage. In Fig. 9, we see the full qutrit simulation (left panel) statistics match fairly well to the RPA (two-level) simulation (right panel), with fluctuations likely caused by finite sampling.
Figure 9: Detection event fraction (DEF) under RPA for thermal model. Distance 3 surface code where all leakage comes from ‘thermal heating’ as discussed in the main text, the same data as in Fig. 1.
Next we consider the second leakage model of interest, with strong CZ leakage, and no thermal heating, as in Sect. III.1.2. As a reminder, here all leakage comes from the diabatic CZ gate, via transitions $|11\rangle\rightarrow\sqrt{1-p}|11\rangle+e^{i\phi_{11,02}}\sqrt{p}|02\rangle$, where the first qubit in the pair (the one that leaks) is always a data qubit. We set the leakage probability to a relatively large value, $p=2.4\times 10^{-3}$.
First we demonstrate in Fig. 10 that the transition phase, $\phi_{11,02}$, has little affect on leakage populations. As such, we can justify setting this phase to 0 in the main text simulations. In contrast however, in Figs. 11, 12 we see that the qubit non-linearity does strongly affect the logical error rate and leakage population. Since the RPA is invariant under changes to the non-linearity (by definition), this can lead to discrepancies in logical error properties (as we saw in Fig. 2).
Figure 10: Leakage populations with varying CZ transition phase. For the distance 3 surface code, with strong CZ leakage (cf. Fig. 2), we plot the two state population for all data qubits, where from left to right the CZ transition phase is $\phi_{11,02}=0,\pi/4,\pi/2$. The CZ transition phase between $|11\rangle,|02\rangle$ is applied during $|11\rangle\rightarrow\sqrt{1-p}|11\rangle+e^{i\phi_{11,02}}\sqrt{p}|02\rangle$. Note, these simulations were preformed without invoking the RPA, taking at least 250k trajectories. The non-linearity is set to $\eta=0.2$GHz.
Figure 11: Logical error probability for different non-linearities. We compare the logical error added by leakage for two values of non-linearity $\eta$, and for the RPA, which is agnostic of this parameter. The leakage model is that from the main text Fig. 2. Data over at least 450k trajectories.
@@ Appendix F Additional RPA simulation data
In Fig. 13 we study the leakage added DEFs. Again, notice that the RPA (right panel) appears to wash out certain coherent effects. In particular, the RPA groups all data in two classes; the four central measure qubits (top four curves), and the four boundary qubits (lower four curves). The asymmetry in the left panel is therefore due to coherent effects not captured by the RPA.
@@
We emphasize here again that this model of strong CZ leakage overestimates coherent effects expected from recent experiments (e.g. scaling_SC_Google_2022 ). Below we show that for more physically chosen parameters, that the RPA does more accurately reproduce the statistics of interest.
@@
Figure 13: Detection event fraction under RPA for CZ leakage. For the distance 3 surface code with strong CZ leakage (cf. Fig. 2), we plot leakage added DEF. The left panel is the full qutrit simulation, and the right panel employs the RPA. Here the non-linearity $\eta=0.2$GHz, and the transition phase $\phi_{11,02}=0$. Data is shown for all measure qubits (see Fig. 7), over at least 450k trajectories.
Appendix G RPA with physical error model
@@
Here we show results pertaining to the repetition/surface codes performance under the RPA, where we use approximately physical noise parameters.
@@
In particular we consider a noise model with two forms of leakage, i) from diabatic CZ gates which introduce transitions $|11\rangle\rightarrow\sqrt{1-p}|11\rangle+e^{i\phi_{11,02}}\sqrt{p}|02\rangle$ with probability $p=4\times 10^{-4}$ and we set $\phi_{11,02}=0$ (see Fig. 10), and ii) a thermal heating model with characteristic time-scale (governing $|1\rangle\rightarrow|2\rangle$) $500\mu$s. For the CZ leakage, we assume the worst case for the code, where the data qubits leak (since measure qubits are reset to $|0\rangle$ each round after measurement). The characteristic leakage time-scale (governing $|2\rangle\rightarrow|1\rangle$ transitions) is $10\mu$s. Additionally the model we use has a $T_{1}$ time (driving $|1\rangle\rightarrow|0\rangle$ transitions) of $20\mu$s, and a dephasing time-scale $T_{\phi}=40\mu$s. The qutrit non-linearity is 200MHz. We discuss the implementation of these models and show the circuits in App. D. Note that each round in our memory experiment takes $1.1\mu$s. Our choice of parameters above is motivated by recent experimental works, such as state-preservation-rc-google ; scaling_SC_Google_2022 ; google-leakage-2016 .
G.1 Repetition code
@@
Overall with these parameters we find a high quantitative agreement between the RPA and the full qutrit simulation. In Fig. 14 we study the extracted logical error rate $\epsilon_{L}$ (see App. E) of the repetition code for distances up to 9. The logical error rate is a key metric in quantum error correction as it determines how long a quantum memory can be stored. We see clearly that the RPA scales as the full simulation does, with no obvious system-size error dependence.
@@
We can additionally consider more ‘microscopic’ properties, such as the distribution of detection events. This is a useful quantity since it is directly related to the noise model implemented. As a reminder, we call the detection event fraction (DEF) the probability a particular stabilizer measurement registers a detection event, in a given round, where a detection event means the stabilizer value changes from one round to the next. In Fig. 15 we observe in the distance 9 repetition code that both with and without the RPA the DEF added by leakage is close to 1.2%.
@@
Figure 14: Logical error rate for repetition code under RPA. Main figure: Here we compare the extracted logical error rates from a $Z$ basis repetition code memory experiment, in simulations of dimension 2 (blue, circle), simulations including leakage (green, diamond), and simulations with leakage and using the RPA (orange, square). Inset: The error rate increase caused by leakage. This is computed by taking the difference of each $D=3$ data point with the corresponding $D=2$ data. Simulation parameters are described in the main text. Data is over at least 200k samples per data point.
@@
Figure 15: Detection event fraction (DEF) of distance 9 repetition code. The probability of a detection event, per stabilizer (measure qubit) per round. Left shows the full qutrit simulation and right invoking the RPA. The long time average for the left plot is 1.18% vs. 1.13% in the right plot with RPA. Data here is over around 225k samples each.
G.2 Surface code
@@
Next we consider a rotated surface code memory experiment at distance 3. In Fig. 16 we see that broadly the logical error probabilities with and without the RPA match well with at most 4% percentage error during the intermediate rounds which can likely be explained to a large extent by sampling errors. The extracted LERs are 0.0248, 0.0283, 0.0284 for the cases without leakage, with leakage, and with leakage using the RPA. In Fig. 17 we study the DEF. Here we see the RPA reproduces the desired results with remarkably high accuracy; in both cases we observe around a 1.15% and 2.05% increase in the number of detection events under leakage, for the boundary and central stabilizers respectively.
@@
Figure 16: Logical error of distance 3 surface code under the RPA. Main figure: The logical error probability for a model with realistic noise parameters, for dimension $D=2$ (blue, circles), and for dimension $D=3$ qutrits (square and diamond). The data of the orange squares corresponds to making the RPA. Inset: The added logical error probability by leakage (i.e. the difference between the two curves and the dimension $D=2$ data). Data is from over 1.5M (0.7M) samples with(out) the RPA. Error bars represent the standard error.
@@
Figure 17: Detection event fraction (DEF) of distance 3 surface code with and without RPA. Left shows the detection event fraction added by leakage, per stabilizer (see legend) per round for a full qutrit simulation. The right is the same but employing the RPA. Note the two curve groupings correspond to the central stabilizers (which undergo 4 CZ gates per round) versus the boundary (flap) stabilizers (2 CZ per round). The long time mean for the former is 2.06% in both cases, and the latter 1.15% (1.16%) with(out) the RPA. The data used is the same as Fig. 16.
Appendix H Classical rate equation
@@
H.1 Incoherent leakage
In the case where leakage is ‘incoherent’ (such as arising form the thermal heating discussed in App. D, and Eq. (25)), we can produce a simple classical model of leakage accumulation. We first assume the 0 and 1 state have equal populations after each round. This is approximately true (up to the effects of noise) for a stabilizer state in the surface code and in bit-flip the repetition code when $X$ flips are performed each round (see Fig. 6). We define the per-round probability for a decay from state $2$ to the computational subspace as $\Gamma_{-}=T_{r}/T_{L}$, where $T_{L}$ is the expected lifetime of state $2$ and $T_{r}$ is the duration of an error correction round. Since the $1\rightarrow 2$ transition time is defined as $T_{h}/2$ (see App. D), the per-round excitation probability to state $2$ is $\Gamma_{+}=\frac{1}{2}\frac{2T_{r}}{T_{h}}=T_{r}/T_{h}$, with $T_{h}\gg T_{L}$ (the factor of $1/2$ in front is by the assumption that $P_{0}\approx P_{1}$, so heating from the computational subspace occurs at half the rate; we do not consider $0\rightarrow 2$ transitions). Letting $P_{c}(k)$ ($P_{2}(k)$) denote the probability of a data qubit being in the computational (leaked) subspace at round $k$, we have the following difference equation:
@@ H.1 Incoherent leakage
H.2 Coherent leakage
In the case leakage is coherent in origin (e.g. via the CZ gate, see Eq. (27)), we find the above model fails to accurately predict leakage populations. We can, however, construct a more detailed classical three-level Markov process to model leakage. Translating the parameters of this model to the thermal leakage model and applying the RPA gives a more accurate match to the coherent leakage populations.
@@ H.2 Coherent leakage
We can fit the model in Eq. (35) to the leakage population data over time. Assuming $\gamma_{01}$ and $\gamma_{12}$ are set by the physical decay times, the model has one unspecified parameter, $\gamma_{21}$, governing the leakage transition $|1\rangle\rightarrow|2\rangle$. The rate extracted by the fit can then be used to construct an effective thermal model, as in the ‘Thermal Approximation’ in Fig. 2A, for which the RPA is an accurate approximation.

### v2

@@ Appendix D Logical error rate
Appendix E Additional STA simulation data
Here we provide supplemental plots and data pertaining to the validation of the Subspace Twirling Approximation, Sect. III.2. First, we show that the detection event statistics relating to the main text ‘thermal heating model’, Fig. 2, are consistent in the model with and without the STA. A detection event is defined as a particular stabilizer value flipping from one round to the next (due to a physical error) chen_exponential_2021 . We compute the detection event fraction (DEF), which is the probability a given measure qubit (stabilizer) registers a detection event, in a given round. It is more instructive to plot the added DEF due to leakage, obtained by considering the difference in DEFs between simulations with and without leakage. In Fig. 10, we see the full qutrit simulation (left panel) statistics match fairly well to the STA (two-level) simulation (right panel), with fluctuations likely caused by finite sampling.
Figure 10: Detection event fraction (DEF) under STA for thermal model. Distance 3 surface code where all leakage comes from ‘thermal heating’ as discussed in the main text, the same data as in Fig. 2.
Next we consider the second leakage model of interest, with strong CZ leakage, and no thermal heating, as in Sect. III.2.2. As a reminder, here all leakage comes from the diabatic CZ gate, via transitions $|11\rangle\rightarrow\sqrt{1-p}|11\rangle+e^{i\phi_{11,02}}\sqrt{p}|02\rangle$, where the first qubit in the pair (the one that leaks) is always a data qubit. We set the leakage probability to a relatively large value, $p=2.4\times 10^{-3}$.
First we demonstrate in Fig. 11 that the transition phase, $\phi_{11,02}$, has little affect on leakage populations. As such, we can justify setting this phase to 0 in the main text simulations. In contrast however, in Figs. 12, 13 we see that the qubit non-linearity does strongly affect the logical error rate and leakage population. Since the STA is invariant under changes to the non-linearity (by definition), this can lead to discrepancies in logical error properties (as we saw in Fig. 3).
Figure 11: Leakage populations with varying CZ transition phase. For the distance 3 surface code, with strong CZ leakage (cf. Fig. 3), we plot the two state population for all data qubits, where from left to right the CZ transition phase is $\phi_{11,02}=0,\pi/4,\pi/2$. The CZ transition phase between $|11\rangle,|02\rangle$ is applied during $|11\rangle\rightarrow\sqrt{1-p}|11\rangle+e^{i\phi_{11,02}}\sqrt{p}|02\rangle$. Note, these simulations were preformed without invoking the STA, taking at least 250k trajectories. The non-linearity is set to $\eta=0.2$GHz.
Figure 12: Logical error probability for different non-linearities. We compare the logical error added by leakage for two values of non-linearity $\eta$, and for the STA, which is agnostic of this parameter. The leakage model is that from the main text Fig. 3. Data over at least 450k trajectories.
@@ Appendix E Additional STA simulation data
In Fig. 14 we study the leakage added DEFs. Again, notice that the STA (right panel) appears to wash out certain coherent effects. In particular, the STA groups all data in two classes; the four central measure qubits (top four curves), and the four boundary qubits (lower four curves). The asymmetry in the left panel is therefore due to coherent effects not captured by the STA.
@@
We emphasize here again that this model of strong CZ leakage overestimates coherent effects expected from recent experiments (e.g. scaling_SC_Google_2022 ). Below we show that for more physically chosen parameters, that the STA does more accurately reproduce the statistics of interest.
@@
Figure 14: Detection event fraction under STA for CZ leakage. For the distance 3 surface code with strong CZ leakage (cf. Fig. 3), we plot leakage added DEF. The left panel is the full qutrit simulation, and the right panel employs the STA. Here the non-linearity $\eta=0.2$GHz, and the transition phase $\phi_{11,02}=0$. Data is shown for all measure qubits (see Fig. 8), over at least 450k trajectories.
Appendix F STA with physical error model
@@
Here we show results pertaining to the repetition/surface codes performance under the STA, where we use approximately physical noise parameters.
@@
In particular we consider a noise model with two forms of leakage, i) from diabatic CZ gates which introduce transitions $|11\rangle\rightarrow\sqrt{1-p}|11\rangle+e^{i\phi_{11,02}}\sqrt{p}|02\rangle$ with probability $p=4\times 10^{-4}$ and we set $\phi_{11,02}=0$ (see Fig. 11), and ii) a thermal heating model with characteristic time-scale (governing $|1\rangle\rightarrow|2\rangle$) $500\mu$s. For the CZ leakage, we assume the worst case for the code, where the data qubits leak (since measure qubits are reset to $|0\rangle$ each round after measurement). The characteristic leakage time-scale (governing $|2\rangle\rightarrow|1\rangle$ transitions) is $10\mu$s. Additionally the model we use has a $T_{1}$ time (driving $|1\rangle\rightarrow|0\rangle$ transitions) of $20\mu$s, and a dephasing time-scale $T_{\phi}=40\mu$s. The qutrit non-linearity is 200MHz. We discuss the implementation of these models and show the circuits in App. C. Note that each round in our memory experiment takes $1.1\mu$s. Our choice of parameters above is motivated by recent experimental works, such as state-preservation-rc-google ; scaling_SC_Google_2022 ; google-leakage-2016 .
F.1 Repetition code
@@
Overall with these parameters we find a high quantitative agreement between the STA and the full qutrit simulation. In Fig. 15 we study the extracted logical error rate $\epsilon_{L}$ (see App. D) of the repetition code for distances up to 9. The logical error rate is a key metric in quantum error correction as it determines how long a quantum memory can be stored. We see clearly that the STA scales as the full simulation does, with no obvious system-size error dependence.
@@
We can additionally consider more ‘microscopic’ properties, such as the distribution of detection events. This is a useful quantity since it is directly related to the noise model implemented. As a reminder, we call the detection event fraction (DEF) the probability a particular stabilizer measurement registers a detection event, in a given round, where a detection event means the stabilizer value changes from one round to the next. In Fig. 16 we observe in the distance 9 repetition code that both with and without the STA the DEF added by leakage is close to 1.2%.
@@
Figure 15: Logical error rate for repetition code under STA. Main figure: Here we compare the extracted logical error rates from a $Z$ basis repetition code memory experiment, in simulations of dimension 2 (blue, circle), simulations including leakage (green, diamond), and simulations with leakage and using the STA (orange, square). Inset: The error rate increase caused by leakage. This is computed by taking the difference of each $D=3$ data point with the corresponding $D=2$ data. Simulation parameters are described in the main text. Data is over at least 200k samples per data point.
@@
Figure 16: Detection event fraction (DEF) of distance 9 repetition code. The probability of a detection event, per stabilizer (measure qubit) per round. Left shows the full qutrit simulation and right invoking the STA. The long time average for the left plot is 1.18% vs. 1.13% in the right plot with STA. Data here is over around 225k samples each.
F.2 Surface code
@@
Next we consider a rotated surface code memory experiment at distance 3. In Fig. 17 we see that broadly the logical error probabilities with and without the STA match well with at most 4% percentage error during the intermediate rounds which can likely be explained to a large extent by sampling errors. The extracted LERs are 0.0248, 0.0283, 0.0284 for the cases without leakage, with leakage, and with leakage using the STA. In Fig. 18 we study the DEF. Here we see the STA reproduces the desired results with remarkably high accuracy; in both cases we observe around a 1.15% and 2.05% increase in the number of detection events under leakage, for the boundary and central stabilizers respectively.
@@
Figure 17: Logical error of distance 3 surface code under the STA. Main figure: The logical error probability for a model with realistic noise parameters, for dimension $D=2$ (blue, circles), and for dimension $D=3$ qutrits (square and diamond). The data of the orange squares corresponds to making the STA. Inset: The added logical error probability by leakage (i.e. the difference between the two curves and the dimension $D=2$ data). Data is from over 1.5M (0.7M) samples with(out) the STA. Error bars represent the standard error.
@@
Figure 18: Detection event fraction (DEF) of distance 3 surface code with and without STA. Left shows the detection event fraction added by leakage, per stabilizer (see legend) per round for a full qutrit simulation. The right is the same but employing the STA. Note the two curve groupings correspond to the central stabilizers (which undergo 4 CZ gates per round) versus the boundary (flap) stabilizers (2 CZ per round). The long time mean for the former is 2.06% in both cases, and the latter 1.15% (1.16%) with(out) the STA. The data used is the same as Fig. 17.
Appendix G Classical rate equation
@@
G.1 Thermal leakage
In the case where leakage arises form the thermal heating discussed in App. C, and Eq. (27)), we can produce a simple classical model of leakage accumulation. We first assume the 0 and 1 state have equal populations after each round. This is approximately true (up to the effects of noise) for a stabilizer state in the surface code and in bit-flip the repetition code when $X$ flips are performed each round (see Fig. 7). We define the per-round probability for a decay from state $2$ to the computational subspace as $\Gamma_{-}=T_{r}/T_{L}$, where $T_{L}$ is the expected lifetime of state $2$ and $T_{r}$ is the duration of an error correction round. Since the $1\rightarrow 2$ transition time is defined as $T_{h}/2$ (see App. C), the per-round excitation probability to state $2$ is $\Gamma_{+}=\frac{1}{2}\frac{2T_{r}}{T_{h}}=T_{r}/T_{h}$, with $T_{h}\gg T_{L}$ (the factor of $1/2$ in front is by the assumption that $P_{0}\approx P_{1}$, so heating from the computational subspace occurs at half the rate; we do not consider $0\rightarrow 2$ transitions). Letting $P_{c}(k)$ ($P_{2}(k)$) denote the probability of a data qubit being in the computational (leaked) subspace at round $k$, we have the following difference equation:
@@ G.1 Thermal leakage
G.2 Coherent leakage
In the case leakage is coherent in origin (e.g. via the CZ gate, see Eq. (29)), we find the above model fails to accurately predict leakage populations. We can, however, construct a more detailed classical three-level Markov process to model leakage. Translating the parameters of this model to the thermal leakage model and applying the STA gives a more accurate match to the coherent leakage populations.
@@ G.2 Coherent leakage
We can fit the model in Eq. (37) to the leakage population data over time. Assuming $\gamma_{01}$ and $\gamma_{12}$ are set by the physical decay times, the model has one unspecified parameter, $\gamma_{21}$, governing the leakage transition $|1\rangle\rightarrow|2\rangle$. The rate extracted by the fit can then be used to construct an effective thermal model, as in the ‘Thermal Approximation’ in Fig. 3A, for which the STA is an accurate approximation.